An excited state in an atom has a life time of 3.0×10⁻⁷ s. Calculate the line width of the emitted photon, when the atom decays spontaneously to its ground state.
1
Expert's answer
2021-10-31T18:12:22-0400

Given:

"\Delta t=3.0*10^{-7}\:\rm s"

The line width is given by

"\Delta \nu=\frac{1}{2\pi\Delta t}=\frac{1}{2\pi*3.0*10^{-7}\:\rm s}=5.3*10^5\:\rm Hz"